OpenStudy (anonymous):

find the derivative of the function w(x)=(24e^x)/(4^x) w'(x)= ??

OpenStudy (anonymous):

\[\frac{dw(x)}{dx}= \frac{\frac{d}{dx}(24*e^x)}{4^x}+\frac{24*e^x}{\frac{ d }{ dx }(4^x)}\] \[\frac{dw(x)}{dx}=\frac{24*(\frac{d}{dx}e^x)}{4^x}+\frac{24*e^x}{\frac{ d }{ dx }(4^x)}\\][\frac{dw(x)}{dx}=\frac{24*e^x}{4^x}+\frac{48*e^x*\ln(2)}{4^x}\]

OpenStudy (anonymous):

Do you need further explanation of the second terms derivative?

OpenStudy (anonymous):

Sorry, that last term is multiplied by a (-1)